An object with a mass of 20 kg is resting above the inclined plane with an angle of 48 degrees above horizontal. What is the normal force?
1
Expert's answer
2020-10-05T10:55:51-0400


Let us draw all the forces acting on the object. From the picture, writing the projections of forces along the axis, perpendicular to the surface, obtain: "N - m g \cos \theta = 0", hence normal force is:

"N = m g \cos \theta = 20 kg \cdot 9.81 \frac{m}{s^2} \cos 48^{\circ} \approx 131.3 N".